OpenStudy (anonymous):

y<1/2x+6, x+3y>+12 how do i graph those inequalities?

OpenStudy (anonymous):

draw the graph of y=1/2x+6 i.e.slope=1/2 and y intercept 6 then the inequality is the one in origin side then draw y=-1/3x+4 the inequality is again in origin side your final answer is the one with common portion..